OpenStudy (esa_pau):

Determine which of the given numbers are roots of the polynomial f(x)= 6x^2+x-1

OpenStudy (esa_pau):

a. 1/2, 1 b.-1/2,2 c. -1/2 d. 1 i know its not B

OpenStudy (vuriffy):

Just plug them in and try.

OpenStudy (vuriffy):

6x^2+x-1 = 0 Let's try -1/2 6(-1/2)^2-1.5 = 0 6(1/4) - 1.5 = 0 1.5 - 1.5 = 0. That means it is a solution.

OpenStudy (vuriffy):

-1/2 - 1 = - 1.5, by the way.

OpenStudy (vuriffy):

0 = 0.

OpenStudy (esa_pau):

alright got it thanks

OpenStudy (vuriffy):

Sweet, great job.

OpenStudy (sshayer):

\[6x^2+3x-2x-1=3x(2x+1)-1(2x+1)=(2x+1)(3x-1)\] 2x+1=0 x=-1/2 3x-1=0 3x=1 x=1/3 from here you can see your option.

OpenStudy (esa_pau):

u lost me

OpenStudy (vuriffy):

The polynomial had two solutions, but your question only gave one of them, which is why I only demonstrated one, whereas @sshayer did both.

OpenStudy (esa_pau):

ahh so it just gave me half an answer in the multiple choice?

OpenStudy (vuriffy):

Technically, yes, it did. But, you only need the one which had -1/2, so it was okay.